OpenStudy (shesolitt):

A cylindrical vase has a diameter of 6 inches. At the bottom of the vase, there are 8 marbles, each of diameter 3 inches. The vase is filled with water up to a height of 10 inches. Which of the following could be used to calculate the volume of water in the vase? π(10in)2(3in) − 1.5(four over threeπ(8in)3) π(3in)2(10in) − 1.5(four over threeπ(8in)3) π(10in)2(3in) − 8(four over threeπ(1.5in)3) π(3in)2(10in) − 8(four over threeπ(1.5in)3) these are the answer choices

Vocaloid (vocaloid):

hint: write an expression to calculate the volume occupied inside the vase

OpenStudy (shesolitt):

i thought it would be A... am i wrong?

Vocaloid (vocaloid):

not quite

Vocaloid (vocaloid):

volume of the vase = (pi)(h)(r^2) = ?

OpenStudy (shesolitt):

so then the answer would be C? im sorry im really bad at geo

Vocaloid (vocaloid):

nope, please answer my question

Vocaloid (vocaloid):

(pi)(h)(r^2) = ?

OpenStudy (shesolitt):

but it says the height of the water in the vase is 10...

Vocaloid (vocaloid):

look carefully at the formula, where does the exponent 2 go?

OpenStudy (shesolitt):

by the radius

Vocaloid (vocaloid):

right, so what is the expression for volume?

OpenStudy (shesolitt):

im just going to go with D.